OpenStudy (anonymous):

I keep getting no solution could I have some help please... Solve the following system of equations. 3x + 3y + 2z = –5 x + y + 3z = –4 2x + 2y + 5z = –7

OpenStudy (lgbasallote):

x = -4 - y - 3z ^substitute to one of the equations 2(-4 - y - 3z) + 2y + 5z = -7 -8 - 2y -6z + 2y + 5z = - 7 -8 - z = -7 z = -8 + 7 z = -1 can you do the rest?

OpenStudy (anonymous):

honestly no I cannot

OpenStudy (lgbasallote):

let's use x = -4 - y - 3z again now sub -1 x = -4 - y -3(-1) x = -4 - y + 3 x = -1 - y got it?

OpenStudy (anonymous):

well why don't you use cramer's rule

OpenStudy (anonymous):

yes thank you very much I understand it now. I really appreciate your help.

OpenStudy (lgbasallote):

great :D